OpenStudy (anonymous):

complete the square on 2x^2-4x+5, leave in vertex form.

OpenStudy (anonymous):

wat we need to do

OpenStudy (anonymous):

here is a snap way vertex is when \(x=-\frac{b}{2a}=-\frac{-4}{2\times 2}=1\) and if \(x=1, y=2\times 1-4\times 1+5=3\) to complete the square therefore, write \[2x^2-4x+5=2(x-1)^2+3\]

OpenStudy (anonymous):

 i think now u can solve furtheer?